OpenStudy (anonymous):

solve the equation. (Use n as an integer constant.) 3 cot^2(x) − 1 = 0 x = (smaller value) x = (larger value)

OpenStudy (ash2326):

\[3 \cot^2 x-1=0\] Add 1 to both the sides \[3 \cot^2 x=1\] Divide by 3 both the sides \[3 \cot^2 x=1\] We get \[\cot^2 x=1/3\] Take root both the sides \[\cot x=\pm\frac{1}{\sqrt 3}\] Do you understand till here

OpenStudy (anonymous):

so far i do, yes.

jhonyy9 (jhonyy9):

so but this not is the end of calcules !!!

OpenStudy (ash2326):

now tell me for what value of x, cot x is \(\frac{1}{\sqrt 3}\) ??

jhonyy9 (jhonyy9):

cotx=sqrt3 /3

jhonyy9 (jhonyy9):

so than x equal what degree ?

jhonyy9 (jhonyy9):

do you know the table of sin,cos,tan,cot for degree 30,45,60 ?

OpenStudy (anonymous):

I am a little familiar with them

OpenStudy (anonymous):

I think i am supposed to be adding 2pi or something.

OpenStudy (anonymous):

my book has an example of another question and its final two answers look like, x=(pi/6)+npi and x =(5pi/6) +npi

OpenStudy (ash2326):

@mirandaqt are you here?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

We had found \[\cot x= \pm \frac{1}{\sqrt 3}\] We know tan x= 1/ cotx so \[\tan x= \pm \sqrt 3\] Now tell for what value of x is tan x= \(\sqrt 3 \)

OpenStudy (anonymous):

i get 60

OpenStudy (anonymous):

yes... 60 degrees.. now, also, going from @ash2326 last post, what other angle gives a tangent of -sqrt3 ?

OpenStudy (anonymous):

-60. but this part is confusing me. my book has an example of another question and its final two answers look like, x=(pi/6)+npi and x =(5pi/6) +npi

OpenStudy (anonymous):

give the positive angle that's coterminal with -60 degrees.